OpenStudy (anonymous):

Add and answer in simplest fractional form. -1/5 + -3/4 = A. -4/9 B. -19/20 C. -4/20

OpenStudy (anonymous):

a

sam (.sam.):

b

OpenStudy (anonymous):

im not sure

sam (.sam.):

\[(-1/5)-(3/4) = -\] Its B

sam (.sam.):

use your calculator

OpenStudy (anonymous):

k

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

You've got to find the lowest common denominator, which is 20 here. Then multiply both fractions by a number that will give you something over 20. Also, both fractions are negative. (4/4)(-1/5) + (5/5)(-3/4) = -(4/20) - (15/20) = -19/20 = Answer B